zarkam21:

? A point at the end of a propeller has a centripetal acceleration of 532.6 m/s^2. If the velocity of the propeller at this point is 35 m/s, how long is the propeller blade?

zarkam21:

1 attachment
zarkam21:

@Shadow

Nnesha:

img008.JPG

sillybilly123:



zarkam21:

so C

zarkam21:

@Nnesha @Vocaloid

Nnesha:

wha formula did you use? and how did you get 1.9m?

zarkam21:

i used the one in the pic

Nnesha:

alright so how did you setup the values ?

zarkam21:

like this

zarkam21:

9.8=35^2/532.6

Nnesha:

\[9.8 \cancel{=}35^2/532.6\]

Nnesha:

\[\frac{ (35)^2 }{ 532.6} \cancel{=} 9.8\]

zarkam21:

0.13

Vocaloid:

a = v^2/r in this case a is the centripetal acceleration given in the problem, v is the velocity of the propeller (also given) and you must solve for r

zarkam21:

532.6=35^2/r

Vocaloid:

awesome, now solve for r

zarkam21:

2.3

Vocaloid:

good, so D